OpenStudy (anonymous):

L x (L+5) = 84 Solve for L

OpenStudy (anonymous):

L x (L+5) = 84 Solve for L

OpenStudy (anonymous):

explain further please

OpenStudy (anonymous):

@Kanwar245

OpenStudy (anonymous):

L^2 + 5L - 84 = 0

OpenStudy (anonymous):

Solve for L

OpenStudy (anonymous):

Im having trouble .

OpenStudy (anonymous):

rewrite it as L^2 + 12L -7L - 84 = 0

OpenStudy (anonymous):

L(L+12)-7(L+12) = 0 (L-7)(L+12) = 0 => L = 7 or L = -12 since L can't be negative, L = 7, therefore the other side is 7 + 5 = 12

OpenStudy (anonymous):

okay I see what you did there. thank you